OpenStudy (anonymous):

Solve the equation. Show your work. 16^3 = 4x

sam (.sam.):

\[16^3 = 4^x??????\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3ln16=xln4 x=6

sam (.sam.):

Don't use ln, \[16^3=4^x\] \[(4^{2})^3=4^x\] \[4^6=4^x\] x=6

OpenStudy (anonymous):

natural logs make the world an easy place my friend; use them, and use them well.